# STIRAP simulations from the 1S to the 2S states for Hydrogen/Antihydrogen atoms

### v2

@@ Abstract
Achieving a high population of antihydrogen/hydrogen atoms in the 2S level is essential for spectroscopy measurements testing similarities between matter and antimatter. We propose and examine the efficiency of applying the STIRAP (Stimulated Raman Adiabatic Passage) process in achieving high population transfer from the 1S to the 2S levels. We utilized a circularly polarized Lyman alpha, Ly-$\alpha$, pulse to couple the 1S state to the 2P state and a microwave pulse to couple the 2P state and the 2S state. We calculate the efficiency of the STIRAP process for transferring the population between the stretched states $(1S_{d},2S_{d})$ as a function of experimental parameters such as Rabi frequencies and pulse durations. We find that a Ly-$\alpha$ pulse with an energy of a few nanojoules could produce nearly perfect transfer at zero detunings for atoms on the laser beam axis. We extended the analysis to a thermal ensemble of atoms, where Doppler detuning affects the velocity distribution of the hydrogen atoms produced in the 2S level. We found that the width of such velocity distribution is controlled by the Rabi frequency. We show that the peak velocity of the hydrogen atoms in the 2S level after STIRAP can be controlled by the Ly-$\alpha$ pulse detuning. The efficiency of STIRAP in transferring population increases at low temperature (Tsimilar-$\sim$1 mK). Finally, we show that a background magnetic field improves the transfer rates between the other trappable states $(1S_{c},2S_{c})$.
@@ I Introduction
Typically, the 1S-2S transition of the hydrogen atom is achieved by using 2 counter-propagating photons of wavelength 243 nm each parthey2011improved ; matveev2013precision . This approach removes the net first order Doppler shift, thus it mostly eliminates the Doppler broadening effects and produces a narrow transition linewidth. Additionally, the hydrogen atoms don’t experience recoil when absorbing the two counter propagating photons. Therefore, this scheme is useful for spectroscopy measurements of the 1S-2S transition. However, when the aim is to produce a large population of hydrogen atoms in the 2S level, it can face some challenges. For instance, the 1S-2S transition is highly nonresonant in the first 243 nm photon of this scheme. Thus, the transition requires high laser intensities parthey2011improved ahmadi2017observation reaching orders of magnitude $10^{7}$ $W\,m^{-2}$ for meaningful transition probabilities. At such high intensities, the rate of photon ionization from the 2S state becomes high rasmussen2017aspects , thus reducing the efficiency of the 1S-2S transfer process. The current efficiency of such a scheme in the literature is about 0.1% rasmussen2017aspects . However, as far as we know, that technique has not been optimized to produce maximum population transfer.
@@
In this work, we simulate the implementation of a STIRAP process to transfer hydrogen or antihydrogen atoms from the 1S to the 2S level. We use the 2P level as the intermediate state, relying on the resonant dipole transitions between the 2P levels and both the initial 1S and the target 2S levels. Two laser pulses are utilized; the Probe, P, pulse, which is a Lyman alpha pulse, Ly-$\alpha$, that couples the 1S and the 2P levels and a Stokes, S, pulse which is a microwave pulse that couples the 2P and the 2S levels. The intensities required for either pulse are of the order $10^{4}$ $W\,m^{-2}$, which is orders of magnitude less than required for excitation using two 243 nm photons. Consequently, the ionization rate, which is proportional to the laser intensity rasmussen2017aspects , is greatly reduced. A possible limitation to applying STIRAP to this system is the high decay rate from the intermediate 2P level to the ground state compared to the almost negligible decay rate into the 2S level. This high decay rate as well as the disparity in the branching ratios are known to reduce the efficiency of the STIRAP process ivanov2005spontaneous . Another possible limitation of applying this approach is that the Doppler shift experienced by the Ly-$\alpha$ pulse is five orders of magnitude higher than that experienced by the microwave pulse connecting the 2S and the 2P levels kramida2010critical . For thermal ensembles of hydrogen atoms at millikelvin temperatures, this shift results in single and two-photon detunings comparable to transition widths of interest. This affects the efficiency of the STIRAP process in transferring populations into the 2S level for higher temperature experiments. On the plus side, the Doppler effect can also be used to generate a population of atoms in the 2S level with a specific range of velocities, Sec. III.2.1. Previous studies have been conducted on the velocity distribution of atomic systems after STIRAP for cooling purposes ivanov2012theory ; ivanov2012robust ; vovk2019features .
@@ II.1 Treatment of STIRAP between the 1S and 2S states of hydrogen
$\vec{E}_{p}(t)=-\mathfrak{Re}\left(E^{0}_{p}e^{-i\left(\omega_{\alpha}+\delta_{p}\right)t}e^{-(2\ln{2})(t-t_{0}^{p})^{2}/\tau^{2}}\hat{\epsilon}_{+}\right)$ (1)
with the unit vector $\hat{\epsilon}_{+}=\frac{-1}{\sqrt{2}}(\hat{x}+i\hat{y})$, $E^{0}_{p}$ is the amplitude of the electric field, the Ly-$\alpha$ transition angular frequency $\omega_{\alpha}=2\pi\times 2.466\times 10^{15}$ rad/s kramida2010critical , $\delta_{p}$ is the detuning in the angular frequency, the FWHM of the intensity duration, $\tau$, was taken to be $20$ ns in all our simulations to be similar to the expected time scale in the proposed HAICU experiments. The pulse also has a Gaussian radial profile with the maximum amplitude of the electric field $E^{0}_{p}$ given by:
@@
where, $w_{0}$ is the waist of the Gaussian pulse, $\mathcal{E}$ is the energy in the pulse, $c$ is the speed of light, and $\epsilon_{0}$ is the permittivity of free space. In many simulations, we used a pulse waist $w_{0}=2$ mm and a total energy of $3$ nJ. We made such choices for the pulse parameters to match those proposed in experiments on trapped hydrogen and antihydrogen atoms for the HAICU project capra2024observation ; fujiwara2023fundamental at TRIUMF (Vancouver, Canada). This pulse produces a Hamiltonian matrix element, as shown in App.A.1, of the form:
@@
In the complete treatment of the STIRAP process, there are 16 possible states in the system; namely 4 hyperfine states for the 1S level, 4x2 states for the 2P (J = 3/2) level, and 4 hyperfine states for the 2S level as illustrated in Fig. 1. However, the circularly polarized pulses, Eq. (1), impose well-known selection rules for a rank one tensor operator georgi2000lie . Due to these selection rules, only 5 of the 2P level states ($F$,$m_{F}$) couple to the S level states, namely the 2P states: $\{(2,2),(2,1),(2,0),(1,1),(1,0)\}$. Thus, resulting in only 13 of the 16 possible states being accessible through the circularly polarized pulses. Thus, the STIRAP Hamiltonian will be a matrix of dimensions $13\times 13$.
@@ II.2 Thermal distribution and recoil effects
We restrict the analysis to the states: $\left\{1S_{d}\otimes\left|k\right>,2P_{a}^{\uparrow}\otimes\left|k+k_{\alpha}\right>,2S_{d}\otimes\left|k+k_{\alpha}\right>\right\}$, where the first part is the internal state of the hydrogen atom and the second part is the center-of-mass momentum state in the direction of the incident photon. The states with the principle quantum number $n=2$ have gained a momentum kick $\hbar k_{\alpha}=2\pi\hbar/(121.6\times 10^{-9}\leavevmode\nobreak\ m)$ from absorbing the Ly-$\alpha$ photon.
@@
The decay transfers population from the state $2P_{a}^{\uparrow}\otimes\left|k+k_{\alpha}\right>$ to the state $1S_{d}\otimes\left|k^{\prime}\right>$, where $k^{\prime}$ could take values in the range $[k:k+2k_{\alpha}]$. The exact value of momentum of the hydrogen atom after emitting a photon depends on the direction of the emitted photon compared to the atom momentum. A crude approximation would be to use the average outcome of multiple decay channels. Since hydrogen atoms are equally likely to emit a photon in the forward and the backward directions, the average effect would be no change in the momentum of hydrogen atoms after decaying to the 1S state, i.e $k^{\prime}=k+k_{\alpha}$. We term this the “averaged approach”.
@@
[…]
@@ II.3 Adding magnetic field effects
$1S_{c}$ $=$ $\left|1,0,0\right>\otimes\big{(}\beta_{1S}\left|1/2,-1/2\right>+\alpha_{1S}\left|-1/2,1/2\right>\big{)}$
@@
$2S_{c}$ $=$ $\left|2,0,0\right>\otimes\big{(}\beta_{2S}\left|1/2,-1/2\right>+\alpha_{2S}\left|-1/2,1/2\right>\big{)}$
@@ III.1 Stationary single-atom $1S_{d}\rightarrow 2S_{d}$ STIRAP
Having made experimentally motivated choices for the parameters of the P pulse: $\Omega^{0}_{p},\tau$ and the pulse waist $w_{0}$, there remain other experimental parameters that could be tuned to produce an efficient STIRAP process, namely $\{\Omega^{0}_{s},\sigma,t_{0}^{p}\}$. To understand the effects of these parameters, we investigate the ideal case in which the atoms are fixed in space at the center of the Ly-$\alpha$ beam. In order to know what values for these parameters would produce an effective STIRAP process, we initialized the hydrogen atom in the $1S_{d}$ state and swept the values of those parameters over the ranges: $\Omega^{0}_{s}\in(0:5]\Omega^{0}_{p}$, $\sigma\in(0:10]\tau$ and $t_{0}^{p}\in(0:2]\sigma$. The optimum values for these parameters were found to be $\Omega^{0}_{s}=\Omega^{0}_{p}$, $\sigma=\tau$. The optimal pulse delay was found to be $t_{0}^{p}=10.5\leavevmode\nobreak\ ns\sim 0.525\tau$, which is smaller than $\sim 0.85\tau$ which is typical in the literature vitanov1997analytic . This difference is attributed to the high decay rate of the 2P level, App. B, as well as the disparity in its branching ratios to the 1S states compared to the negligible decays into the 2S states. To test that, we ran separate simulations for smaller decay rates $\sim 0.01\Gamma_{0}$ and equal branching ratios and both lead to more population transfer to the 2S level at larger pulses’ separations when keeping the other pulses’ parameters fixed. The case of equal branching ratios had better agreement with the $\sim 0.85\tau$ in the literature RevModPhys.89.015006 . However, for the decay rates and the branching ratios of the 2P states of the hydrogen atom, App. B, a pulses’ separation of $t_{0}^{p}=10.5\leavevmode\nobreak\ ns\sim 0.525\tau$ gave the best population transfer into the 2S level. Unless otherwise indicated, we use the pulses’ parameters listed in this paragraph for the rest of the simulations in this work.
@@
Figure 3: The final probability to be in the $2S_{d}$ state as a function of the P pulse detuning $\delta_{p}$ for the Rabi frequencies of the S and P pulses $\Omega^{0}_{p}=\Omega^{0}_{s}=\{1,2,4,8,16\}\times\Omega_{0}$, with $\Omega_{0}=4.915\times 10^{8}$ rad/sec. The simulations were run while keeping the microwave pulse detuning $\delta_{s}$ as well as the pulses duration $\tau=\sigma=20$ ns fixed; thus, changing the Ly-A pulse energy. The transfer rate to the 2S level drops for $\delta_{p}$ outside the range $[-\Omega^{0}_{p}:\Omega^{0}_{p}]$.
@@ A.1 Stationary atoms STIRAP Hamiltonian
$\Omega_{p}$ $=-\frac{eE^{0}_{p}}{\hbar\sqrt{2}}\left<1|x-iy|2\right>$ $=-eE^{0}_{p}\left<1|r_{-}|2\right>/\hbar$
$\Omega_{s}$ $=-\frac{eE^{0}_{s}}{\hbar\sqrt{2}}\left<3|x-iy|2\right>$ $=-eE^{0}_{s}\left<3|r_{-}|2\right>/\hbar$ (24)
@@ Appendix B Decay channels and rates
$\Gamma_{1}(\left|2,2\right>\rightarrow\left|1\right>)$ $=$ $\Gamma_{0}$
$\Gamma_{2}(\left|2,1\right>\rightarrow\left|1\right>)$ $=$ $\Gamma_{0}/2$
$\Gamma_{3}(\left|2,1\right>\rightarrow\left|+\right>)$ $=$ $\Gamma_{0}/2$
$\Gamma_{4}(\left|2,0\right>\rightarrow\left|1\right>)$ $=$ $3\displaystyle 2\Gamma_{0}/$ 3
$\Gamma_{5}(\left|2,0\right>\rightarrow\left|+\right>)$ $=$ $\Gamma_{0}/6$
$\Gamma_{6}(\left|2,0\right>\rightarrow\left|-1\right>)$ $=$ $\Gamma_{0}/6$
$\Gamma_{7}(\left|1,1\right>\rightarrow\left|1\right>)$ $=$ $\Gamma_{0}/6$
$\Gamma_{8}(\left|1,1\right>\rightarrow\left|+\right>)$ $=$ $\Gamma_{0}/6$
$\Gamma_{9}(\left|1,1\right>\rightarrow\left|-\right>)$ $=$ $3\displaystyle 2\Gamma_{0}/$ 3
$\Gamma_{10}(\left|1,0\right>\rightarrow\left|1\right>)$ $=$ $3\displaystyle 2\Gamma_{0}/$ 3
$\Gamma_{11}(\left|1,0\right>\rightarrow\left|-\right>)$ $=$ $\Gamma_{0}/6$
$\Gamma_{12}(\left|1,0\right>\rightarrow\left|-1\right>)$ $=$ $\Gamma_{0}/6$ (29)

### v3

@@ Abstract
[…]
@@ I Introduction
Typically, the 1S-2S transition of the hydrogen atom is achieved by using 2 counter-propagating photons of wavelength 243 nm each parthey2011improved ; matveev2013precision . This approach removes the net first order Doppler shift, thus it mostly eliminates the Doppler broadening effects and produces a narrow transition linewidth. Additionally, the hydrogen atoms don’t experience recoil when absorbing the two counter propagating photons. Therefore, this scheme is useful for spectroscopy measurements of the 1S-2S transition. However, when the aim is to produce a large population of hydrogen atoms in the 2S level, it can face some challenges. For instance, the 1S-2S transition is highly nonresonant in the first 243 nm photon of this scheme. Thus, the transition requires high laser intensities and energies parthey2011improved ahmadi2017observation reaching orders of magnitude $10^{7}$ $W\,m^{-2}$ and $10^{-3}$ J, respectively, for meaningful transition probabilities. At such high intensities, the rate of photo-ionization from the 2S state becomes high rasmussen2017aspects , thus reducing the efficiency of the 1S-2S transfer process. The current efficiency of such a scheme in the literature is about 0.1% rasmussen2017aspects . However, as far as we know, that technique has not been optimized to produce maximum population transfer. Other studies have investigated population transfer through the continuum carroll1992coherent ; nakajima1994population ; yatsenko1997population . However, in yatsenko1997population it was shown that when such technique is applied to hydrogen, the ionization rates are comparable to those seen in the 243 nm photons scheme.
@@
In this work, we simulate the implementation of a STIRAP process to transfer hydrogen or antihydrogen atoms from the 1S to the 2S level. We use the 2P level as the intermediate state, relying on the resonant dipole transitions between the 2P levels and both the initial 1S and the target 2S levels. Two radiation pulses are utilized; the Probe, P, pulse, which is a Lyman alpha pulse, Ly-$\alpha$, that couples the 1S and the 2P levels and a Stokes, S, pulse which is a microwave pulse that couples the 2P and the 2S levels. The required Ly-$\alpha$ pulse has intensities of $\sim 10^{4}$ W m-2 and energies of $\sim 10^{-9}$ J, several orders of magnitude less than those used in the 243 nm two-photon excitation. Consequently, the ionization probability becomes negligible (of order $10^{-8}$ of the excited atoms). A possible limitation to applying STIRAP to this system is the high decay rate from the intermediate 2P level to the ground state compared to the almost negligible decay rate into the 2S level. This high decay rate as well as the disparity in the branching ratios are known to reduce the efficiency of the STIRAP process ivanov2005spontaneous . Another possible limitation of applying this approach is that the Doppler shift experienced by the Ly-$\alpha$ pulse is five orders of magnitude higher than that experienced by the microwave pulse connecting the 2S and the 2P levels kramida2010critical . For thermal ensembles of hydrogen atoms at millikelvin temperatures, this shift results in single and two-photon detunings comparable to transition widths of interest. This affects the efficiency of the STIRAP process in transferring populations into the 2S level for higher temperature experiments. On the plus side, the Doppler effect can also be used to generate a population of atoms in the 2S level with a specific range of velocities, Sec. III.2.1. Previous studies have been conducted on the velocity distribution of atomic systems after STIRAP for cooling purposes ivanov2012theory ; ivanov2012robust ; vovk2019features .
@@ II.1 Treatment of STIRAP between the 1S and 2S states of hydrogen
[…]
@@
where, $w_{0}$ is the waist of the Gaussian pulse, $\mathcal{E}$ is the energy in the pulse, $c$ is the speed of light, and $\epsilon_{0}$ is the permittivity of free space. In many simulations, we used a pulse waist $w_{0}=2$ mm and a total energy of $3$ nJ. We made such choices for the pulse parameters to match those proposed in experiments on trapped hydrogen and antihydrogen atoms for the HAICU project fujiwara2020demonstration ; capra2024observation ; fujiwara2023fundamental at TRIUMF (Vancouver, Canada). This pulse produces a Hamiltonian matrix element, as shown in App.A.1, of the form:
@@
In the complete treatment of the STIRAP process, there are 16 possible states in the system; namely 4 hyperfine states for the 1S level, 4x2 states for the 2P (J = 3/2) level, and 4 hyperfine states for the 2S level as illustrated in Fig. 1. However, the circularly polarized pulses, Eq. (1), impose well-known selection rules for a rank one tensor operator georgi2000lie . Due to these selection rules, only 5 of the 2P level states ($Fm_{F}$) couple to the S level states, namely the 2P states: $\{(2,2),(2,1),(2,0),(1,1),(1,0)\}$. Thus, resulting in only 13 of the 16 possible states being accessible through the circularly polarized pulses. Thus, the STIRAP Hamiltonian will be a matrix of dimensions $13\times 13$.
@@ II.2 Thermal distribution and recoil effects
[…]
@@
The decay transfers population from the state $2P_{a}^{\uparrow}\otimes\left|k+k_{\alpha}\right$> to the state $1S_{d}\otimes\left|k^{\prime}\right$>, where $k^{\prime}$ could take values in the range $k:k+2k_{\alpha}]$. The exact value of momentum of the hydrogen atom after emitting a photon depends on the direction of the emitted photon compared to the atom momentum. A crude approximation would be to use the average outcome of multiple decay channels. Since hydrogen atoms are equally likely to emit a photon in the forward and the backward directions, the average effect would be no change in the momentum of hydrogen atoms after decaying to the 1S state, i.e $k^{\prime}=k+k_{\alpha}$. We term this the “averaged approach”.
@@
$\xi(k^{\prime})=\frac{3}{8k_{\alpha}}\int_{k^{\prime}-\delta k/2}^{k^{\prime}+\delta k/2}\left(1+f(\bar{k})^{2}\right)d\bar{k}$ (7)
@@ II.3 Adding magnetic field effects
$2P_{a}^{\downarrow}$ $=$ $\left|2,1,1\right>\otimes\left|1/2,-1/2\right$> (9)
[…]
@@ III.1 Stationary single-atom $1S_{d}\rightarrow 2S_{d}$ STIRAP
Having made experimentally motivated choices for the parameters of the P pulse: $\Omega^{0}_{p},\tau$ and the pulse waist $w_{0}$, there remain other experimental parameters that could be tuned to produce an efficient STIRAP process, namely $\{\Omega^{0}_{s},\sigma,t_{0}^{p}\}$. To understand the effects of these parameters, we investigate the ideal case in which the atoms are fixed in space at the center of the Ly-$\alpha$ beam. In order to know what values for these parameters would produce an effective STIRAP process, we initialized the hydrogen atom in the $1S_{d}$ state and swept the values of those parameters over the ranges: $\Omega^{0}_{s}\in(0:5]\Omega^{0}_{p}$, $\sigma\in(0:10]\tau$ and $t_{0}^{p}\in(0:2]\sigma$. The optimum values for these parameters were found to be $\Omega^{0}_{s}=\Omega^{0}_{p}$, $\sigma=\tau$. The optimal pulse delay was found to be $t_{0}^{p}=10.5\penalty 10000\ ns\sim 0.525\tau$, which is smaller than $\sim 0.85\tau$ which is typical in the literature vitanov1997analytic . This difference is attributed to the high decay rate of the 2P level, App. B, as well as the disparity in its branching ratios to the 1S states compared to the negligible decays into the 2S states. To test that, we ran separate simulations for smaller decay rates $\sim 0.01\Gamma_{0}$ and equal branching ratios and both lead to more population transfer to the 2S level at larger pulses’ separations when keeping the other pulses’ parameters fixed. The case of equal branching ratios had better agreement with the $\sim 0.85\tau$ in the literature RevModPhys.89.015006 . However, for the decay rates and the branching ratios of the 2P states of the hydrogen atom, App. B, a pulses’ separation of $t_{0}^{p}=10.5\penalty 10000\ ns\sim 0.525\tau$ gave the best population transfer into the 2S level. We examined implementing microwave pulses proportional to $e^{-t^{4}/\sigma^{4}}$ and obtained results similar to a Gaussian pulse. Unless otherwise indicated, we use the pulses’ parameters listed in this paragraph for the rest of the simulations in this work.
@@
Figure 3: The final probability to be in the $2S_{d}$ state as a function of the P pulse detuning $\delta_{p}$ for the Rabi frequencies of the S and P pulses $\Omega^{0}_{p}=\Omega^{0}_{s}=\{1,2,4,8,16\}\times\Omega_{0}$, with $\Omega_{0}=4.915\times 10^{8}$ rad/sec. The simulations were run while keeping the microwave pulse detuning $\delta_{s}$ as well as the pulses duration $\tau=\sigma=20$ ns fixed; thus, changing the Ly-A pulse energy. The transfer rate to the 2S level drops for $\delta_{p}$ outside the range $\Omega^{0}_{p}:\Omega^{0}_{p}]$.
@@ IV Conclusion
As discussed in Sec. III.1 we experimented with different possible parameters for the microwave pulse’s configuration, and the ones used in the simulations were the optimal ones found. The Ly-$\alpha$ pulse was restricted to a Gaussian pulse for experimental purposes. Further STIRAP optimization techniques such as optimal pulse control and the Stimulated Raman Exact Passage (STIREP) stefanatos2021optimal ; laforgue2022optimal ; liu2023optimal could be future research directions if needed.
@@ A.1 Stationary atoms STIRAP Hamiltonian
$\Omega_{p}$ $=-\frac{eE^{0}_{p}}{\hbar\sqrt{2}}\left<1|x-iy|2\right$> $=-eE^{0}_{p}\left<1|r_{-}|2\right$$\hbar$
$\Omega_{s}$ $=-\frac{eE^{0}_{s}}{\hbar\sqrt{2}}\left<3|x-iy|2\right$> $=-eE^{0}_{s}\left<3|r_{-}|2\right$$\hbar$ (24)
@@ Appendix B Decay channels and rates
[…]